OpenStudy (anonymous):

Please help! In the interval [0,pi) what are the angles for cos x = square root 2 / 2 ?

OpenStudy (anonymous):

only one in that interval is 45 degrees, or if you are working in radian \(\frac{\pi}{4}\)

OpenStudy (anonymous):

look at unit circle on the last page of the attached cheat sheet cosine is the first coordinate. you will see that where the first coordinate \(\frac{\sqrt{2}}{2}\) the angle is \(\frac{\pi}{4}\)

OpenStudy (anonymous):

Wow thank you so much! That cheat sheet helps a lot!